A copper vessel of mass 135 kg contains 6.75 kg of water at a temperature of 25C. Find the heat roquited to warm the vessel and water to 90°C. Take specific heat of copper = 0 406 JAg K and specific heat of water 4.187 kkg K.


1
Expert's answer
2021-04-08T07:34:51-0400

The heat required to increase the temperature of the vessel:


"Q_v=c_vm_v\Delta t,\\
Q_w=c_wm_w\Delta t,\\
Q=Q_v+Q_v=\Delta t(c_vm_v+c_wm_w)=\\
=(90-25)(135·0.406+4187·6.75)=1840609\text{ J}."